Fulkerson Manufacturing wishes to maintain a sustainable growth rate of 9 percent a year, a debt-equity ratio of 41, and a dividend payout ratio of 32 percent. The ratio of total assets to sales is constant at 1.34 What profit margin must the frm achieve in order to meet its growth rate goal? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, .g.32.16 Profa margn
Answer #1
we have to first compute the ROE
we know that-
Sustainable growth rate = ROE* retention ratio/(1- ROE* retention ratio)
retention ratio = 1- dividend pay out ratio = 1-32% 68.00%
therefore, we have-
9% = =ROE*68%/(1-ROE*68%)
=0.09- 0.68*.09*ROE = 0.68ROE
=0.09 = 0.7412*ROE
ROE = 12.14%
Using DUPONT formula we have
ROE = Profit margin * Asset turnover ratio*equity multiplier
12.14% = Profit margin * (1/1.34)*(1.41)
Profit margin = =12.14%/((1/1.34)*1.41)
Profit margin = 11.54%